OpenStudy (anonymous):

14/16 x 6/32 divided by 3/4

OpenStudy (anonymous):

first divide 14 by sixteen and get what

OpenStudy (anonymous):

0.875?

OpenStudy (igreen):

Hold on, is it \(\sf (\dfrac{14}{16} \times \dfrac{6}{32}) \div \dfrac{3}{4}\) or is it \(\sf \dfrac{14}{16} \times (\dfrac{6}{32} \div \dfrac{3}{4})\)

OpenStudy (anonymous):

right and then divide 6 into thirty two

OpenStudy (anonymous):

@iGreen there are no parenthesis. If I had a file to attach I would send it to you.

OpenStudy (anonymous):

@dilphill31 0.1875

OpenStudy (anonymous):

and i think u would multiply them answers

OpenStudy (anonymous):

ok one sec.

OpenStudy (anonymous):

hang on let me get one of my friends @rickygore126 @dakid88

OpenStudy (anonymous):

it's a very large number

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think i messed up when i told u to multiply them sorry

OpenStudy (anonymous):

it's fine

OpenStudy (dakid88):

first u combine the fractions to get a common denominator

OpenStudy (anonymous):

ok so the common denominator would be 32 right?

OpenStudy (anonymous):

yes and then u do what

OpenStudy (dakid88):

u will combine 14/16 and 6/32 and the common denominator will be 2

OpenStudy (anonymous):

and then you multiply the top by two

OpenStudy (anonymous):

???? I thought the common denominator was 32 because 16 and 4 can both go into it

OpenStudy (mathmale):

I'd suggest doing some preliminary planning before coming up with "answers." 14/16 x 6/32 divided by 3/4 could be written as \[14/16 * 6/32 .divided. by. 3/4\rightarrow \frac{ \frac{ 14 }{ 16 }*\frac{ 6 }{ 32 } }{ \frac{ 3 }{ 4 }}\]

OpenStudy (dakid88):

yes @rickygore126 i did not mean to put that sorry

OpenStudy (mathmale):

Please do not use " x " to denote multiplication. This is pre-algebra or algebra, in which we use " x " to represent an unknown quantity. Use " * " to denote multiplication. Secondly: note that NO LCD is needed here. Can anyone explain why that is so?

OpenStudy (dakid88):

but the answer will beee.........

OpenStudy (anonymous):

the common denominator is 32 but you would have to multiply the numbers on the bottom to get it and what do you multiply 16 by to get 32

OpenStudy (dakid88):

2

OpenStudy (mathmale):

Once again: There is NO need for a common denominator here. Talking about finding an LCD will only confuse the issue.

OpenStudy (anonymous):

you mulyiply 16 by 2

OpenStudy (dakid88):

exactly

OpenStudy (anonymous):

@mathmale I already know there is no "LCD" I'm not looking for a least common denominator.

OpenStudy (anonymous):

right so you would multiply 14 by two and get what

OpenStudy (anonymous):

28

OpenStudy (mathmale):

\[\frac{ \frac{ 14 }{ 16 }*\frac{ 6 }{ 32 } }{ \frac{ 3 }{ 4 }}\]

OpenStudy (anonymous):

then if u can simplify 28/32

OpenStudy (dakid88):

and tht will give u 7/32

OpenStudy (mathmale):

May I ask what your goal is? What's the purpose of going through all these steps?

OpenStudy (anonymous):

no @dakidd88 because we still have to do some more stuff

OpenStudy (anonymous):

@mathmale to get the answer, obviously

OpenStudy (anonymous):

to answer the problem @mathmale

OpenStudy (anonymous):

let's get back to work @rickygore126

OpenStudy (anonymous):

@mathmale another goal of mine is to eat pizza for lunch

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

Your 14/16 x 6/32 divided by 3/4 translates into \[\frac{ \frac{ 14 }{ 16 }*\frac{ 6 }{ 32 } }{ \frac{ 3 }{ 4 }}\]

OpenStudy (anonymous):

so we have 6/32 and 7/32 so far

OpenStudy (mathmale):

No multiplication is needed here to simplify. Instead, cancel wherever possible.

OpenStudy (anonymous):

@mathmale yes I know because you've sent it 400 times

OpenStudy (anonymous):

to simplify it divide 28 into 32 and what is your new fraction

OpenStudy (anonymous):

if I divide it won't I get a decimal?

OpenStudy (anonymous):

yes and then you would convert that decimal into a fraction and get 7/8

OpenStudy (anonymous):

oh, ok I see . so after we get 7/8 what do you do?

OpenStudy (anonymous):

you set it up and multiply straight across with the other number and get what

OpenStudy (anonymous):

multiple it by 7/32, 6/32, or both?

OpenStudy (anonymous):

multiply 7/8 by 6/32

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it's a decimal answer, so do I convert that to a fraction?

OpenStudy (anonymous):

right and the fraction is 21/128

OpenStudy (mathmale):

May I suggest you summarize what you have accomplished and what you still have to do to complete this problem? Simply by cancelling, I can derive the final answer in a matter of seconds. There is no point to converting these fractions to decimal fractions (unless, of course, you can understand decimal fractions better than proper fractions).

OpenStudy (anonymous):

ok. is that the final answer, or am I leaving something out?

OpenStudy (anonymous):

no that is not your final answer get that answer and we have to divide it by 3/4 which i will show you how in a minute when u get the answer

OpenStudy (anonymous):

@mathmale all you're doing is confusing me because I am trying to do this problem with @dilphill31 and @dakid88 and you keep sending the same message.

OpenStudy (anonymous):

@dilphill31 not you

OpenStudy (anonymous):

i know and dakidd left

OpenStudy (anonymous):

alright have u got he answer

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

ok that is fine

OpenStudy (anonymous):

0.013671875

OpenStudy (anonymous):

and then I need to convert it

OpenStudy (anonymous):

when u convert the 42/256 u get 21/128

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and now the final step

OpenStudy (anonymous):

let's set up the problem and it will be what

OpenStudy (anonymous):

well we have 6/32, 7/32, and 21/128 but I don't know how to set t up

OpenStudy (anonymous):

forget the other fractions when u set it up it will look like this 21/128 divided by 3/4

OpenStudy (anonymous):

oh, true. one sec let me solve it real quick

OpenStudy (anonymous):

didn't we already divide these number though?

OpenStudy (anonymous):

unless you did and i didn't know but you have to keep the first number the same flip the sign to multiplcation and then flip it to make 4/3 and multiply straight across and get what

OpenStudy (anonymous):

I got 84/384 but I don't think that is correct

OpenStudy (anonymous):

hang on let me work it out

OpenStudy (anonymous):

that is right and can we convert it

OpenStudy (anonymous):

do we simplify it down

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got 7/32

OpenStudy (anonymous):

if all that hard work we just did pays off you are correct

OpenStudy (anonymous):

i am 95% sure we done that right and if we got it wrong sorry but i tried the best i could

OpenStudy (anonymous):

Oh my gosh. Thank you so much for staying an hour and helping me with this. you have no idea how much that helped.

OpenStudy (anonymous):

you are welcome